# Draft: May 24, 2025

### v1

@@ Abstract
The demand of finite raw materials will keep increasing as they fuel modern society. Simultaneously, solutions for stopping carbon emissions in the short term are not available, thus making the net zero target extremely challenging to achieve at scale. The circular economy (CE) paradigm is gaining attention as a solution to address climate change and the uncertainties of supplies of critical materials. Hence, in this paper, we introduce CiRL, a deep reinforcement learning (DRL) library of environments focused on the circularity of both solid and fluid materials. The integration of DRL into the design of material circularity is possible thanks to the formalism of thermodynamical material networks, which is underpinned by compartmental dynamical thermodynamics. Along with the focus on circularity, this library has three more features: the new CE-oriented environments are in the state-space form, which is typically used in dynamical systems analysis and control designs; it is based on a state-of-the-art Python library of DRL algorithms, namely, Stable-Baselines3; and it is developed in Google Colaboratory to be accessible to researchers from different disciplines and backgrounds as is often the case for circular economy researchers and engineers. CiRL is publicly available [^1] https://github.com/ciroresearch/CiRL.
Keywords: Waste sorting, air pollution control, nature-based solutions, circular robotics, circular intelligence.
@@ 1 Introduction
While modern societies provide customized and effective services and products to individuals, they rely on strategic materials whose supply is not guaranteed in the long-term (say, in fifty years) because of the finiteness of the natural reserves, the prohibitive costs for further extraction, or the unreliability of the countries from which they are imported (Nygaard, 2023; European Commission, 2023). Another open issue of modern societies is the generation of waste that ends-up in water bodies (Suryawan et al., 2024) and urban areas (Pietrelli et al., 2025; Ballatore et al., 2022). In addition, carbon dioxide emissions resulting from combustion in energy, heating, and aerospace systems continue to increase since today there is no immediate solution to stop them (NASA, ). Hence, the net-zero target requested by several governments to prevent a climate event horizon is extremely challenging to reach (Oxford Net Zero, ; European Commission, a; UK Government, ; Office of the Federal Chief Sustainability Officer, ). The coexistence of these problems motivates the rise of the paradigm of circular economy (CE), in which solid and fluid materials are kept in use for as long as possible through reusing, repairing, refurbishing, remanufacturing, and also through reducing the material demand itself (Ellen MacArthur Foundation, ; Liaros, 2021; Ossio et al., 2023; Potting et al., 2017; Pajunen and Holuszko, 2022).
Deep reinforcement learning (DRL) is a very promising control approach since it enhances classical control techniques by leveraging the latest advances in deep learning. Although not yet performing for highly dynamic real systems, DRL has the potential to yield very adaptive and reliable controllers if combined with the stability property developed in control theory (Buşoniu et al., 2018; Han et al., 2020). The interest for DRL is demonstrated by the proliferation of libraries and physical simulators created in the last four years (Fujita et al., 2021; Weng et al., 2022; Seno and Imai, 2022; Tarasov et al., 2023; Serrano-Muñoz et al., 2023; Gallouédec et al., 2021; Scheikl et al., 2023; Suarez, 2024; Raffin et al., 2021; Kaup et al., 2024).
Despite the above-mentioned benefits of a transition to a circular economy and the potential of DRL for effective decision making, the link between CE and DRL remains very vague in the literature mainly because of the lack of clarity on how to measure the circularity of a system. For example, Halter et al. (2025) found 798 circularity indicators used in the literature. In other words, it is unanswered the question: “How the recent advances in DRL can increase circularity?”
@@
    where $\mathcal{N}$ is a network of thermodynamic compartments that exchange and process a target material, e.g., plastic, gold, carbon dioxide (Zocco et al., 2024).
  • It provides an open-source library of DRL environments to solve (1), namely, CiRL, leveraging the state-of-the-art DRL algorithm implementations of Stable-Baselines3 (Raffin et al., 2021). In addition, CiRL is user-friendly to be accessible to the circular economy community and its environments are in the state-space form to highlight the parallelism with classical control theory, and hence, to reach the classical control community.
@@
The rest of the paper is organized as follows: Section 2 covers the related work, then Section 3 provides the preliminary definitions and concepts necessary to proceed; subsequently, Section 4 details the features of CiRL, Section 5 explains the environments implemented in CiRL so far, while Section 6 gives suggestions to users and developers of CiRL; finally, Section 7 concludes by discussing the limitations and future work.
@@ 2.1 Libraries for Deep Reinforcement Learning
Several DRL libraries have been proposed with different features and applications. One of the first DRL libraries was KerasRL (Plappert, 2016) proposed in 2016 and it is based on the deep learning framework Keras. One of the most popular and extensive libraries is SB3 (Raffin et al., 2021) and was proposed in 2021. Another library proposed in 2021 was ChainerRL (Fujita et al., 2021), which was based on the deep learning framework Chainer. Two libraries proposed in 2022 are Tianshou (Weng et al., 2022) and d3rlpy (Seno and Imai, 2022). The former supports both offline and online training through a unified interface, while the latter focuses on offline learning. Another library focused on offline settings is CORL (Tarasov et al., 2023), which was proposed in 2023. CORL considers also the offline-to-online option. A library supporting a particularly vast set of environment interfaces is skrl because it supports, along with the popular Farama Gymnasium interface (Towers et al., 2024), the NVIDIA Isaac Gym, Isaac Orbit, and Omniverse Isaac Gym environments (Serrano-Muñoz et al., 2023). The libraries mentioned so far do not target specific domains of applications. In contrast, FinRL focuses on quantitative finance (Liu et al., 2020), Panda-Gym focuses on the Franka Emika Panda robot (Gallouédec et al., 2021), while LapGym focuses on robot-assisted laparoscopic surgery (Scheikl et al., 2023). A library proposed in 2024 is PufferLib, whose aim is to facilitate the combination of the algorithms available in libraries, e.g, SB3, and environment simulators (Suarez, 2024). An extensive list of libraries is available in (Wang, ), while a comparison of selected ones is provided by (Nouwou Mindom et al., 2023).
With respect to existing libraries, CiRL: is largely developed through Python notebooks similarly to FinRL, it follows the Open AI Gym/Farama Gymnasium interface similarly to SB3, ChainerRL, and d3rlpy, and it relies on SB3 for the training algorithms; CiRL is the first library aiming to optimize the circularity of materials, namely, $\lambda(\mathcal{N};t)$ (Zocco and Malvezzi, ).
@@ 2.2 Circular Economy
The paradigm of circular economy is gaining interest as a solution to reduce material supply uncertainties (Nygaard, 2023) and waste generation (Luttenberger, 2020) by closing the loop of material flows, e.g., the output materials of a process can enter another process as secondary raw material. Circularity extends the material life-cycle and reduces the demand of finite raw materials, especially those imported from uncertain natural reserves or unreliable areas of the world (Distefano et al., 2024).
At the foundations of circularity are the so called “Rs”, that indicate the practices to be adopted in order to increase circularity. Examples of these practices are the reduction of material demand, the reuse of materials and product components, the repair of parts, and recycling (Potting et al., 2017). The “Rs” are usually ordered from the least to the most energy intensive and the aim is to extend the life-cycle of products and materials at their highest value possible (thus, recycling is one of the last options). The adoption of the “Rs” is being investigated in different sectors, e.g., in healthcare (van Straten et al., 2021), in automotive (Prochatzki et al., 2023), in aerospace (Dias et al., 2022), in food (Liaros, 2021), in fashion (Peleg Mizrachi and Tal, 2022), in electronics (Pajunen and Holuszko, 2022), and in construction (Ossio et al., 2023).
@@
where $\mathcal{N}$ is a network of thermodynamic compartments designed to maximize circularity $\lambda(\mathcal{N};t)$ (Zocco and Malvezzi, ; Zocco et al., 2024). The next section will explain $\mathcal{N}$ and $\lambda(\mathcal{N};t)$ in details.
@@ 3 Preliminaries
Definition 1 (Bondy and Murty (1976))
@@
which transport, store, use, and transform a target material, namely, $\beta$. Each compartment is indicated by a control surface (Moran et al., 2010) and is modeled using dynamical systems derived from a mass balance and/or at least one of the laws of thermodynamics (Haddad, 2019).
Specifically, $\mathcal{N}=\mathcal{R}\cup\mathcal{T}$, where $\mathcal{R}\subseteq\mathcal{N}$ is the subset of compartments $c^{k}_{i,j}$ that store, transform, or use the target material, while $\mathcal{T}\subset\mathcal{N}$ is the subset of compartments $c^{k}_{i,j}$ that move the target material between the compartments belonging to $\mathcal{R}\subseteq\mathcal{N}$. A net $\mathcal{N}$ is associated with its compartmental diagraph $M(\mathcal{N})$, which is a digraph whose nodes are the compartments $c^{k}_{i,j}\in\mathcal{R}$ and whose arcs are the compartments $c^{k}_{i,j}\in\mathcal{T}$. For node-compartments $c^{k}_{i,j}\in\mathcal{R}$ it holds that $i=j=k$, whereas for arc-compartments $c^{k}_{i,j}\in\mathcal{T}$ it holds that $i\neq j$ because an arc moves the material from the node-compartment $c^{i}_{i,i}$ to the node-compartment $c^{j}_{j,j}$. The orientation of an arc is given by the direction of the material flow. The superscript $k$ is the identifier of each compartment. The superscripts $k_{\text{v}}$ and $k_{\text{a}}$ in (3) are the $k$-th node and the $k$-th arc, respectively, while $n_{\text{c}}$ and $n_{\text{v}}$ are the total number of compartments and nodes, respectively. Since $n_{\text{a}}$ is the total number of arcs, it holds that $n_{\text{c}}=n_{\text{v}}+n_{\text{a}}$ (Zocco et al., 2023; Zocco, 2024; Zocco et al., 2024).
@@
Definition 3 (Zocco and Malvezzi )
@@
Definition 4 (Zocco and Malvezzi )
The instantaneous circularity $\lambda(\mathcal{N};t)\in(-\infty,0]$ is defined as
@@ 4 Features of CiRL
    Focused on material circularity: To the best of our knowledge, CiRL is the first RL library whose goal is to optimize $\lambda(\mathcal{N};t)$ (4). The intersection between RL and CE is possible thanks to the formalism of TMNs (Definition 2) (Zocco et al., 2023, 2024; Zocco, 2024; Zocco et al., 2025b; Zocco and Malvezzi, ), where one or more thermodynamic compartments are controlled by RL agents that aim to maximize $\lambda(\mathcal{N};t)$.
@@
    Developed in Google Colaboratory: CiRL is mainly a collection of Python notebooks written and executed on Google Colaboratory (aka Colab) (Google Colab, ), and hence, it is more user-friendly than libraries developed with plain programming scripts, e.g., Stable-Baselines3 (SB3) (Raffin et al., 2021), Gymnasium (Towers et al., 2024), and TorchRL (Bou et al., 2024). The extensive use of Python notebooks is in common with the FinRL library (Liu et al., 2020). This makes reinforcement learning more accessible to circular economy researchers and engineers whose programming is not the main skill. Since Google maintains the hardware and drivers of Colab, the user needs only to focus on high-level Python programming. To save the training results and models, Google Drive is the default storage location in CiRL. Colab has a free plan suitable for small-scale experiments; subscription and pay-as-you-go plans are both available if more computing resources are needed, e.g., a more powerful GPU.
@@
    Based on SB3: SB3 is a state-of-the-art library of reinforcement learning algorithms actively maintained and updated (Raffin et al., 2021). CiRL takes the RL algorithms directly from SB3 and proposes a first collection of environments for the circulation of solids and fluid materials. Specifically, the thermodynamic compartments of (3) for the circularity of solid materials currently implemented in CiRL are a robot for waste sorting, a truck for transportation, and an incinerator for waste removal and energy generation (Zocco et al., 2024); the only fluid currently addressed in CiRL is the atmospheric carbon dioxide because it is one of the main causes of climate instability (NASA, ). Microalgae systems are the thermodynamic compartments considered in CiRL for carbon removal (Bernard et al., 2016; Zocco et al., 2025a).
@@
    Environments in state-space form: CiRL leverages the robot environments available in Gymnasium Robotics (Towers et al., 2024), but it also implements new environments specific for circular economy designs, e.g., an incinerator, an algal cultivation. The new environments are written in the state-space representation, which is the standard form used in systems and control theory for the analysis of systems and design of controllers (Goodwin et al., 2001). Hence, systems and control researchers and engineers can easily interpret CiRL to design RL controllers.
5 Compartments
In this section, we describe the open-source environments currently created in CiRL. The environments and the tested SB3 algorithms are summarized in Table 1 along with their roles as network compartments. The RL algorithms are: the A2C (Mnih et al., 2016), the augmented random search (ARS) (Mania et al., 2018), the deep deterministic policy gradient (DDPG) (Lillicrap et al., 2015), the proximal policy optimization (PPO) (Schulman et al., 2017), the soft actor-critic (SAC) (Haarnoja et al., 2018), the truncated quantile critics (TQC) (Kuznetsov et al., 2020), and the twin delayed DDPG (TD3) (Fujimoto et al., 2018). The combination of one of these algorithms with the hindsight experience replay (HER) (Andrychowicz et al., 2017) is indicated as “algorithm-HER”, e.g., “DDPG-HER” in the case of DDPG with HER.

### v2

# CiRL: Open-Source Environments for Reinforcement Learning in Circular Economy and Net Zero

@@ Abstract
The demand of finite raw materials will keep increasing as they fuel modern society. Simultaneously, solutions for stopping carbon emissions in the short term are not available, thus making the net zero target extremely challenging to achieve at scale. The circular economy (CE) paradigm is gaining attention as a solution to address climate change and the uncertainties of supplies of critical materials. Hence, in this paper, we introduce CiRL, a deep reinforcement learning (DRL) library of environments focused on the circularity control of both solid and fluid materials. The integration of DRL into the design of material circularity is possible thanks to the formalism of thermodynamical material networks, which is underpinned by compartmental dynamical thermodynamics. Along with the focus on circularity, this library has three more features: the new CE-oriented environments are in the state-space form, which is typically used in dynamical systems analysis and control design; it is based on a state-of-the-art Python library of DRL algorithms, namely, Stable-Baselines3; and it is developed in Google Colaboratory to be accessible to researchers from different disciplines and backgrounds as is often the case for circular economy researchers and engineers. CiRL is intended to be a tool to generate AI-driven actions for optimizing the circularity of supply-recovery chains and to be combined with human-driven decisions derived from material flow analysis (MFA) studies. CiRL is publicly available [^1] https://github.com/ciroresearch/CiRL.
keywords:
Waste sorting , air pollution control , nature-based solutions , circular robotics , circular intelligence.
[…]
organization=Department of Information Engineering and Mathematics, University of Siena, addressline=Via Roma 56, city=Siena, postcode=53100, country=Italy
@@ 1 Introduction
While modern societies provide customized and effective services and products to individuals, they rely on strategic materials whose supply is not guaranteed in the long-term (say, in fifty years) because of the finiteness of the natural reserves, the prohibitive costs for further extraction, or the unreliability of the countries from which they are imported (nygaard2023geopolitical; CRM-EU). Another open issue of modern societies is the generation of waste that ends-up in water bodies (suryawan2024reshaping) and urban areas (pietrelli2025characterization; ballatore2022city). In addition, carbon dioxide emissions resulting from combustion in energy, heating, and aerospace systems continue to increase since today there is no immediate solution to stop them (NASA-ClimateChange). Hence, the net-zero target requested by several governments to prevent a climate event horizon is extremely challenging to reach (NZ-DefUniOx; EU-NZ; UK-NZ; US-NZ). The coexistence of these problems motivates the rise of the paradigm of circular economy (CE), in which solid and fluid materials are kept in use for as long as possible through reusing, repairing, refurbishing, remanufacturing, and also through reducing the material demand itself (EllenMacArthurFound; liaros2021circular; ossio2023circular; potting2017circular; pajunen2022circular).
Deep reinforcement learning (DRL) is a very promising control approach since it enhances classical control techniques by leveraging the latest advances in deep learning. Although not yet performing for highly dynamic real systems, DRL has the potential to yield very adaptive and reliable controllers if combined with the stability property developed in control theory (bucsoniu2018reinforcement; han2020actor). However, challenges remain open such as the sample inefficiency of learning algorithms (zheng2024sample) and the formulation of effective reward functions to solve real-world problems (feng2025deep; lanzaro2025evaluating). The interest for DRL is demonstrated by the proliferation of libraries and physical simulators created in the last four years (fujita2021chainerrl; weng2022tianshou; seno2022d3rlpy; tarasov2023corl; serrano2023skrl; gallouedec2021panda; scheikl2023lapgym; suarez2024pufferlib; raffin2021stable; kaup2024review).
Despite the above-mentioned benefits of a transition to a circular economy and the potential of DRL for effective decision making, the link between CE and DRL remains, to the best of our knowledge, still vague in the literature mainly because of the lack of clarity on how to measure the circularity of a system. For example, halter2025techno found 798 circularity indicators used in the literature. In other words, it is unanswered the question: “How the recent advances in DRL can increase circularity?”
@@
    where $\mathcal{N}$ is a network of thermodynamic compartments that exchange and process a target material, e.g., plastic, gold, carbon dioxide (zocco2024unification).
  • 2.
    It provides an open-source library of DRL environments to solve (1), namely, CiRL, leveraging the state-of-the-art DRL algorithm implementations of Stable-Baselines3 (raffin2021stable). In addition, CiRL is user-friendly to be accessible to the circular economy community and its environments are in the state-space form to highlight the parallelism with classical control theory, and hence, to reach the classical control community.
@@
The rest of the paper is organized as follows: Section 2 covers the related work, then Section 3 provides the preliminary definitions and concepts necessary to proceed; subsequently, Section 4 details the features of CiRL, Section 5 explains the environments implemented in CiRL so far, while Section 6 gives suggestions to users and developers of CiRL. Finally, Section 7 discusses the applications of CiRL, Section 8 discusses the limitations and future work, and Section 9 gives the conclusion.
@@ 2.1 Libraries for Deep Reinforcement Learning
Several DRL libraries have been proposed with different features and applications. in this section, we mention some of the libraries proposed from 2016 to 2024 following a chronological order. One of the first DRL libraries was KerasRL (plappert2016kerasrl) proposed in 2016 and it is based on the deep learning framework Keras. One of the most popular and extensive libraries is SB3 (raffin2021stable) and was proposed in 2021. Another library proposed in 2021 was ChainerRL (fujita2021chainerrl), which was based on the deep learning framework Chainer. RL algorithms can work either off-line or on-line; the former train the agents using datasets of historical data of observations, actions, and rewards; thus, the dataset defines the environment in the off-line setting. In contrast, on-line methods require that the training data are generated while the agent is learning. CiRL follows the mechanics of the latter. Two libraries proposed in 2022 are Tianshou (weng2022tianshou) and d3rlpy (seno2022d3rlpy). The former supports both off-line and on-line training through a unified interface, while the latter focuses on off-line learning. Another library focused on off-line settings is CORL (tarasov2023corl), which was proposed in 2023. CORL considers also the offline-to-online option. A library supporting a particularly vast set of environment interfaces is skrl because it supports, along with the popular Farama Gymnasium interface (towers2024gymnasium), the NVIDIA Isaac Gym, Isaac Orbit, and Omniverse Isaac Gym environments (serrano2023skrl). To date, Isaac Gym is no longer under development, all development and features are in IsaacSim. Isaac Orbit is open source so that contributions can be made by the community; Isaac Orbit now evolves as Isaac Lab. Our library is compatible with the Farama Gymnasium interface only. The libraries mentioned so far do not target specific domains of applications. In contrast, FinRL focuses on quantitative finance (liu2020finrl), Panda-Gym focuses on the Franka Emika Panda robot (gallouedec2021panda), while LapGym focuses on robot-assisted laparoscopic surgery (scheikl2023lapgym). A library proposed in 2024 is PufferLib, whose aim is to facilitate the combination of the algorithms available in libraries, e.g, SB3, and environment simulators (suarez2024pufferlib). An extensive list of libraries is available in (listOfRLlibraries), while a comparison of selected ones is provided by (nouwou2023comparison).
With respect to existing libraries, CiRL: is largely developed through Python notebooks similarly to FinRL, it follows the Open AI Gym/Farama Gymnasium interface similarly to SB3, ChainerRL, and d3rlpy, and it relies on SB3 for the training algorithms; CiRL is the first library aiming to optimize the circularity of materials, namely, $\lambda(\mathcal{N};t)$ (zocco2024circular). What makes CiRL unique is that, to date, is the only library that supports decision making in circular systems designs. Furthermore, it is the only library of environments written in the state-space form; this is to stimulate analysis from and discussions with control theorists, whose control techniques differ from RL. At a high-level, CiRL helps with the European effort to develop digital solutions for a circular economy as promoted, for example, by the DICE Network+ in the United Kingdom (DICENetwork+) and by the DICE Lab in central Europe (DiCE-Lab).
@@ 2.2 Circular Economy
The paradigm of circular economy is gaining interest as a solution to reduce material supply uncertainties (nygaard2023geopolitical) and waste generation (luttenberger2020waste) by closing the loop of material flows, e.g., the output materials of a process can enter another process as secondary raw material. Circularity extends the material life-cycle and reduces the demand of finite raw materials, especially those imported from uncertain natural reserves or unreliable areas of the world (distefano2024material).
At the foundations of circularity are the so called “Rs”, that indicate the practices to be adopted in order to increase circularity. Examples of these practices are the reduction of material demand, the reuse of materials and product components, the repair of parts, and recycling (potting2017circular). The “Rs” are usually ordered from the least to the most energy intensive and the aim is to extend the life-cycle of products and materials at their highest value possible (thus, recycling is one of the last options). The adoption of the “Rs” is being investigated in different sectors, e.g., in healthcare (van2021circular), in automotive (prochatzki2023critical), in aerospace (dias2022possibilities), in food (liaros2021circular), in fashion (peleg2022regulation), in electronics (pajunen2022circular), and in construction (ossio2023circular).
@@
where $\mathcal{N}$ is a network of thermodynamic compartments designed to maximize circularity $\lambda(\mathcal{N};t)$ (zocco2024circular; zocco2024unification). The next section will explain $\mathcal{N}$ and $\lambda(\mathcal{N};t)$ in details.
@@ 3 Preliminaries
Definition 1 (bondy1976graph)
@@
which transport, store, use, and transform a target material, namely, $\beta$. Each compartment is indicated by a control surface (moran2010fundamentals) and is modeled using dynamical systems derived from a mass balance and/or at least one of the laws of thermodynamics (haddad2019dynamical).
Specifically, $\mathcal{N}=\mathcal{R}\cup\mathcal{T}$, where $\mathcal{R}\subseteq\mathcal{N}$ is the subset of compartments $c^{k}_{i,j}$ that store, transform, or use the target material, while $\mathcal{T}\subset\mathcal{N}$ is the subset of compartments $c^{k}_{i,j}$ that move the target material between the compartments belonging to $\mathcal{R}\subseteq\mathcal{N}$. A net $\mathcal{N}$ is associated with its compartmental diagraph $M(\mathcal{N})$, which is a digraph whose nodes are the compartments $c^{k}_{i,j}\in\mathcal{R}$ and whose arcs are the compartments $c^{k}_{i,j}\in\mathcal{T}$. For node-compartments $c^{k}_{i,j}\in\mathcal{R}$ it holds that $i=j=k$, whereas for arc-compartments $c^{k}_{i,j}\in\mathcal{T}$ it holds that $i\neq j$ because an arc moves the material from the node-compartment $c^{i}_{i,i}$ to the node-compartment $c^{j}_{j,j}$. The orientation of an arc is given by the direction of the material flow. The superscript $k$ is the identifier of each compartment. The superscripts $k_{\text{v}}$ and $k_{\text{a}}$ in (3) are the $k$-th node and the $k$-th arc, respectively, while $n_{\text{c}}$ and $n_{\text{v}}$ are the total number of compartments and nodes, respectively. Since $n_{\text{a}}$ is the total number of arcs, it holds that $n_{\text{c}}=n_{\text{v}}+n_{\text{a}}$ (zocco2023thermodynamical; zocco2022circularity; zocco2024unification).
@@
Definition 3 (zocco2024circular)
@@
Definition 4 (zocco2024circular)
The instantaneous circularity $\lambda(\mathcal{N};t)$ is defined as
@@
Remark: The circularity $\lambda(\mathcal{N};t)$ takes into account the mass transfers taking place in a system, but only those that are finite-time sustainable as defined in Definition 3. We derived the notion of circularity (4) by formulating the Ellen MacArthur Foundation’s definition of circular economy (EllenMacArthurFound) in terms of fundamental physics principles, i.e., in terms of mass transfers. The explicit dependence on time indicates that circularity is a time-varying quantity since it is directly affected by the dynamics of the mass transfers in the system. The negative sign before the parentheses in the right-hand side of the equality in (4) was added to state the design for systems circularity as a maximization problem. Indeed, with the minus, reducing the extraction of finite resources, e.g., gold or copper, yields a reduction of the quantity inside the parentheses, and hence, an increase of $\lambda(\mathcal{N};t)$. A good measure of circularity must increase with reductions of raw material mining coherently with the Ellen MacArthur Foundation’s definition (EllenMacArthurFound).
@@ 4 Features of CiRL
    Focused on material circularity: To the best of our knowledge, CiRL is the first RL library whose goal is to optimize $\lambda(\mathcal{N};t)$ (4). The intersection between RL and CE is possible thanks to the formalism of TMNs (Definition 2) (zocco2023thermodynamical; zocco2024unification; zocco2022circularity; zocco2024synchronized; zocco2024circular), where one or more thermodynamic compartments are controlled by RL agents that aim to maximize $\lambda(\mathcal{N};t)$.
@@
    Developed in Google Colaboratory: CiRL is mainly a collection of Python notebooks written and executed on Google Colaboratory (aka Colab) (ColabWebpage), and hence, it is more user-friendly than libraries developed with plain programming scripts, e.g., Stable-Baselines3 (SB3) (raffin2021stable), Gymnasium (towers2024gymnasium), and TorchRL (boutorchrl). The extensive use of Python notebooks is in common with the FinRL library (liu2020finrl). This makes reinforcement learning more accessible to circular economy researchers and engineers whose programming is not the main skill. Since Google maintains the hardware and drivers of Colab, the user needs only to focus on high-level Python programming. To save the training results and models, Google Drive is the default storage location in CiRL. Colab has a free plan suitable for small-scale experiments and the NVIDIA Tesla T4 is the GPU available for free; subscription and pay-as-you-go plans are both available if more computing resources are needed, e.g., a more powerful GPU.
@@
    Based on SB3: SB3 is a state-of-the-art library of reinforcement learning algorithms actively maintained and updated (raffin2021stable). CiRL takes the RL algorithms directly from SB3 and proposes a first collection of environments for the circulation of solids and fluid materials. Specifically, the thermodynamic compartments of (3) for the circularity of solid materials currently implemented in CiRL are a robot for waste sorting, a truck for transportation, and an incinerator for waste removal and energy generation (zocco2024unification); the only fluid currently addressed in CiRL is the atmospheric carbon dioxide because it is one of the main causes of climate instability (NASA-ClimateChange). Microalgae systems are the thermodynamic compartments considered in CiRL for carbon removal (bernard2016modelling; zocco2025circular).
@@
    Environments in state-space form: CiRL leverages the robot environments available in Gymnasium Robotics (towers2024gymnasium), but it also implements new environments specific for circular economy designs, e.g., an incinerator, an algal cultivation. The new environments are written in the state-space representation, which is the standard form used in systems and control theory for the analysis of systems and design of controllers (goodwin2001control). Hence, systems and control researchers and engineers can easily interpret CiRL to design RL controllers.
5 Compartments and Compartmental Networks
In this section, we describe the open-source environments currently created in CiRL. The environments and the tested SB3 algorithms are summarized in Table 1 along with their roles as network compartments. The RL algorithms are: the A2C (mnih2016asynchronous), the augmented random search (ARS) (NEURIPS2018_7634ea65), the deep deterministic policy gradient (DDPG) (lillicrap2019continuouscontroldeepreinforcement), the proximal policy optimization (PPO) (schulman2017proximal), the soft actor-critic (SAC) (haarnoja2018soft), the truncated quantile critics (TQC) (kuznetsov2020controlling), and the twin delayed DDPG (TD3) (fujimoto2018addressing). The combination of one of these algorithms with the hindsight experience replay (HER) (andrychowicz2017hindsight) is indicated as “algorithm-HER”, e.g., “DDPG-HER” in the case of DDPG with HER.